OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. sin 2x + sin x - 2

OpenStudy (anonymous):

@jim_thompson5910 Can you help me understand these?

OpenStudy (anonymous):

sin(2x) = 2sin(x)cos(x)

OpenStudy (anonymous):

Is the last term sin(x-2), or sin(x)-2?

OpenStudy (anonymous):

sin ^2x + sin x - 2 I think it means sin(x). It just has a space there.

jimthompson5910 (jim_thompson5910):

hint: let z = sin(x)

jimthompson5910 (jim_thompson5910):

z^2 = sin^2(x)

jimthompson5910 (jim_thompson5910):

so, sin^2(x) + sin(x) - 2 turns into z^2 + z - 2

jimthompson5910 (jim_thompson5910):

hopefully you see how to go from here

OpenStudy (anonymous):

(sin(x)+2)(sin(x)-1)? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

Awesome! I've got a few more I don't quite understand..

jimthompson5910 (jim_thompson5910):

ok I'll help with 2 more

OpenStudy (anonymous):

Find sin θ if cot θ = - 4 and cos θ < 0. This one is like the one we did earlier, but a little different. I have no idea what it is asking for.

jimthompson5910 (jim_thompson5910):

what is cot in terms of opp, adj, hyp?

OpenStudy (anonymous):

adj/opp

jimthompson5910 (jim_thompson5910):

so cot(theta) = -4 and cot(theta) = adj/opp because -4 = -4/1, we know that adj = -4 opp = 1

jimthompson5910 (jim_thompson5910):

Now draw a right triangle with that info and use it to find the hypotenuse

OpenStudy (anonymous):

sqrt17

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now you can find sin(theta)

OpenStudy (anonymous):

sqrt17/17?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Awesome. Hang on. I working the next one.

OpenStudy (anonymous):

*I'm